1. Given powdered sodium sulfate, how will you prepare 100mL of a 1M solution of sodium sulfate? Your solvent is 0.1M acetate Buffer.
2. Given a 1M solution of sodium chloride, how will you prepare 1000uL of 0.2M solution of sodium chloride.
3. Given a 100 mg/mL protein solution, how will you prepare 20uL of a 70 mg/mL protein solution?
4. What is lysozyme and where are they usually found?
5. What is the role of X-crystallography in structure identification?
Answer #1

Answer 1

1 M solution has 1 mole in 1000 ml solvent

Required volume = 100 ml

So, moles required = 0.1 mole

Molar mass of sodium sulfate = 142.04 g/mol

Mass of sodium sulfate required for solution

= 0.1 mol*142.04 g/mol

= 14.2 g

So, to prepare required solution, take 14.2 g of sodium sulfate and dissolve it in 50 ml of 0.1 M acetate buffer to form clear solution in a beaker. Pour this solution into a volumetric flask of 100 ml. Rinse the beaker with 10-15 ml of 0.1 M acetate buffer and pour the rinsing contents in volumetric flask. Make up the volume to 100 ml by adding acetate buffer till the etched mark.

Answer 2

Given concentration (C1) = 1.0 M

Diluted concentration (C2) = 0.2 M

Diluted volume (V2) = 1000 uL

Required volume of C1 (V1) = to be calculated

Applying dilution formula

C1*V1 = C2*V2

V1 = C2*V2/C1

= (0.2 M* 1000 uL)/(1 M)

= 200 uL

So, for preparing 1000 uL of 0.2 M NaCl solution, measure 200 uL of 1 M solution and add 800 uL of solvent.

Answer 3

C1 = 100 mg/ml

C2 = 70 mg/ml

V1 = to calculate

V2 = 20 uL

Applying dilution formula

V1 = C2*V2/C1

= (70 mg/ml * 20 uL)/(100 mg/ml)

= 1.40 uL

So, , for preparing 20 uL of 70 mg/ml protein solution, measure 1.40 uL of 100 mg/ml solution and add 18.6 uL of solvent.

Answer 5

Lysozyme are hydrolytic enzymes. These hydrolyze the peptide bonds of bacterial cell walls by cleaving bonds between N-acetyl galactosamine and N-acetyl muramic acid.

These are found in human tears, saliva, sweat etc.

Answer 5

X-crystallography involves diffraction of incident X-rays by the crystal whose structure is to be determined. The diffraction pattern is analyzed to determine the three dimensional crystal structure of the sample.